OpenStudy (anonymous):

Factor the polynomial completely. x^3+x^2+2x-4 Using synthetic division.

OpenStudy (anonymous):

I'm here lol xD ; )

OpenStudy (anonymous):

whats synthetic division ? can't I use normal factorization ! O_o

OpenStudy (anonymous):

Its like long division, but faster and simpler.

OpenStudy (anonymous):

If I'm allowed Normal Factorization Ok I get it ! Different Region ...Different Terms ..It's Fast but not on Open Study ; )

OpenStudy (anonymous):

\[x^3 + x^2 + 2x -4 = 0 \]That's the Equation Right ? Now When I put 1for x ! I get \[L.H.S =R.H.S\]

OpenStudy (anonymous):

So It Implies 1 is a root for the Equation ...

OpenStudy (anonymous):

Then I Divide \[\frac{(x^3 + x^2 + 2x - 4)}{x - 1} \]

OpenStudy (anonymous):

I just checked and wow! How did I miss that!? I swear I checked one, but thanks! Now I can finish it. :)

OpenStudy (anonymous):

__________________________ x - 1 | x^3 + x^2 + 2x -4 |x^2 +x + 4 x^3 - x^2 ___________ 2x^2 + 2x 2x^2 - 2x _____________ + 4x - 4 4x - 4 _________ 0 There You Go YOU Have a Equation Again \[ x^2 + x +4=0\] Solve it To Get the other two Roots

OpenStudy (anonymous):

\[x = \frac{ -1 \pm \sqrt{ 1 -4}}{2}\] \[x = \frac{-1 \pm i \sqrt{3} }{2}\] Which Gives Us Complex Roots or Imaginary Roots Thus Roots are \[1,w,w^2\]

OpenStudy (anonymous):

Did you get it ? : )